the z-scores that bound the middle 41% under the standard normal curve are z=-a and z=a. find the value of a
Answer #1

Given P(-a<z<a) = 0.41

To find a

from standard normal table P(z<0.54) = 0.7054

P(Z<-0.54) =0.2946

P(-0.54 < z< 0.54) = P(Z<0.54) - P(z<-0.54) = 0.7054-0.2946 = 0.4108 = 41%

Hence a = 0.54